If the Diels-Alder reaction is classified as a [4this the pre lab questions for The Diels - Alder Cycloaddition Reactio expirment please answer all the question
Answer #1

1. The [4+2]-cycloaddition of a conjugated diene and a dienophile (an alkene or alkyne), an electrocyclic reaction that involves the 4 π-electrons of the diene and 2 π-electrons of the dienophile

a) 4+2 since 4 pi electrons of the alkene and 2 free electrons are taking part in the reaction

b) 2+2 : since both the pi electrons are taking place in the reaction

2. Density = mass/Volume

0.8 = 0.14/V (converted the mg to g)

V = 0.14/0.8 = 0.175 mL

3. Limiting reagent always depends on the number of moles of the substance which is used in the reaction.

First convert the amount of substances given into its moles by dividing it with their molar mass. and now see which is the least amount that will be the limiting reagent

4. No it does not undergo diel alder reaction because the diene should be conjugated

The Diels–Alder reaction is an organic chemical reaction (specifically, a [4+2] cycloaddition) between a conjugated diene and a substituted alkene, commonly termed the dienophile, to form a substituted cyclohexene system